OpenStudy (anonymous):

Help! I give out medals! What is the x-intercept of the line given by the equation x minus four y equals twenty? The x-intercept is 0. The x-intercept is –4. The x-intercept is 5. The x-intercept is 20.

OpenStudy (anonymous):

? Help!

OpenStudy (akashdeepdeb):

The equation is: x - 4y = 20 The graph may look something like this.  To find the x-intercept we know that the y value MUST be 0, do you know why? Because the x-intercept is the point where the line touches the x-axis, and when the line toughes the x-axis, the y-value of the point is zero [No Rise]. So put y=0 And find the value of x. Understood this? :)

rvc (rvc):

20 @hotgurl72

OpenStudy (anonymous):

OK. But he says it's 20. ?

OpenStudy (anonymous):

Which is it, 20 or 0?

OpenStudy (akashdeepdeb):

@hotgurl72 Did you get my explanation? :)

OpenStudy (anonymous):

YEs. I guess.

OpenStudy (akashdeepdeb):

I told you to put the value of y=0. And then just find 'x'! :D Because if the equation given to you is, x-4y = 20 Then you can easily get the value of x! Try now, what do you think the answer is? :)

OpenStudy (anonymous):

What is the y-intercept of the line? Graph shows a line with negative slope running through points four comma zero and zero comma six The y-intercept is 0. The y-intercept is 4. The y-intercept is 6. The y-intercept is Negative six.

OpenStudy (akashdeepdeb):

The answer as @rvc said is 20. This is because, x-4y = 20 ALWAYS REMEMBER: To find x-intercept y=0 To find y-intercept x=0 Here we have to find the x-intercept. So y=0 x - 4(0) = 20 x - 0 = 20 x = 20 Thus, that is the x-intercept of the line [x=20]. Did you get WHY this is the case?

OpenStudy (anonymous):

Help?

OpenStudy (akashdeepdeb):

Now try finding out the y-intercept of your new question from the hints I have given you. :)

OpenStudy (anonymous):

So x is 20, not 0?

OpenStudy (anonymous):

Just say yes or no please.

OpenStudy (anonymous):

It's 20.

rvc (rvc):

yes!

rvc (rvc):

@hotgurl72

OpenStudy (anonymous):

OK, I have another question: If Joelle multiplied 792 by a positive integer and came up with a perfect square as her answer, then what is the smallest integer she could have multiplied 792 by?

rvc (rvc):

wats792*2

rvc (rvc):

@hotgurl72

OpenStudy (akashdeepdeb):

@hotgurl72 We are not allowed to give out answers on this website, and for a very good reason at that! If I just give you the answer it may help you for the problem Set you are doing but it is not helpful AT ALL in the long run (which is what matters). Please go through the OpenStudy Code-Of-Conduct http://www.openstudy.com/code-of-conduct Everyone will help you with the question here! :)